OpenStudy (anonymous):

Sum the series to n terms - 1.4.7 + 4.7.10 + 7.10.13 +...........................to n terms.

OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

\[a_n=(3n+1)(3n+4)(3n+7)\]

OpenStudy (anonymous):

\[n=0,1,2,...\]

OpenStudy (anonymous):

i think there will be (3n-2) in place of 3n+7

OpenStudy (anonymous):

n starts from 0 ... check again :)

OpenStudy (anonymous):

but n is natural no.

OpenStudy (anonymous):

it doesn't matter to use natural or start from 0 ha?

OpenStudy (anonymous):

because the result is same

OpenStudy (anonymous):

these series are applicable only when n is natural.

OpenStudy (anonymous):

ok we turn it to natural...\[a_n=(3n-2)(3n+1)(3n+4)\]\[n=1,2,3,...\]

OpenStudy (anonymous):

now?

OpenStudy (anonymous):

one way (uglyiest) is to multiply all of brackets and using \(\sum k\) , \(\sum k^2\) and \(\sum k^3\)

OpenStudy (anonymous):

but there are nice methods...

OpenStudy (anonymous):

??

OpenStudy (anonymous):

guys what do u think?

OpenStudy (anonymous):

after simplifying u have\[a_n=27n^3+27n^2-18n-8\]

OpenStudy (anonymous):

i have solved like this- \[a_{n} = 27n^{3}+27n^{2}-18n-8\] \[S_{n}=\sum{a_{n}}\] \[S_{n}=27\sum{n^{3}}+27\sum{n^{2}}-18\sum{n}-8n\]

OpenStudy (anonymous):

exactly

mathslover (mathslover):

did i interrupt mukushla? I think I did 'answer snipping' that is a hard one for me .. sorry you may continue though I promise I will interrupt at last

OpenStudy (anonymous):

\[\sum k^3=(\frac{n(n+1)}{2})^2\]\[\sum k^2=\frac{n(n+1)(2n+1)}{6}\]\[\sum k=\frac{n(n+1)}{2}\]

OpenStudy (anonymous):

i know all these values, continue.........

OpenStudy (anonymous):

@mathslover np ... :)

OpenStudy (anonymous):

\[S_n=\frac{n}{4}(27n^3+90n^2+45n-50)\]this is what u will got after simplifying according to god of wolf

OpenStudy (anonymous):

can it be simplified more

OpenStudy (anonymous):

???

OpenStudy (anonymous):

it must divisible by 4 so yeah i think it can...let me ask wolf

OpenStudy (anonymous):

it is the ans.

OpenStudy (anonymous):

see its expanded form and match with yours.

OpenStudy (anonymous):

so its ok

OpenStudy (anonymous):

yes it is correct

OpenStudy (anonymous):

how will you simplify it further?

OpenStudy (anonymous):

The above series can be written in the form: tn = (1+3n)^3 -9(1+3n), where n=1,2,3...

OpenStudy (anonymous):

thanks i got it.